# Unavoidable Canonical Nonlinearity Induced by Gaussian Measures Discretization

### v2

@@ II Concept and Derivation
Eq.(II) certainly exhibits that the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure.
This result establishes a clear geometric interpretation of the UCN: The cost to project continuous Gaussian onto a discrete configurational support is given by the cumulative information-geometric distance for expected parallel translations within the Gaussian statistical manifold. These contributions are invisible to existing KL-based comparisons restricted to discrete statistical manifolds, but are naturally captured by the present transport-information-geometric framework.
@@
In the previous sections, we demonstrated that the discrepancy between continuous Gaussian distribution and its discretized counterpart, measured by the 2-Wasserstein distance with a quadratic cost aligned with Fisher metric, converges in the vanishing discretization limit $d\to 0$ to the expectation of the KL divergence between infinitesimally translated continous Gaussian. Although these derivations was shown for Gaussian families, the underlying mathematical structure indicates its broader applicability. We here briefly address the conditions where such a correspondence can be generalized.
@@
where $f$ corresponds to the number of effective degrees of freedom relevant to the description of the system interested (and, subsequently, to the transport problem). Then we consider the problem to measure the difference between function $P\in S$ and its discretized counterpart $P^{\prime}$, based on $W_{2}$ and KL-divergence as discussed for Gaussian families. We assume that, in these coordinates, the Fisher metric (or its block-diagonal component associated with the transported degrees of freedom) is given by $f\times f$ positive-definite matrix of $\Omega$, which is taken as constant. More generally, our statement rely on the locally-constant character of $\Omega$ under inifitesimal variations in $\xi$. Then we define the quadratic cost function for the 2-Wasserstein distance $W_{2}\left(P:P^{\prime}\right)$ as
@@
$\rho\left(\delta\xi\right)=\textrm{Unif}\left[-\frac{h\left(d\right)}{2},\frac{h\left(d\right)}{2}\right]^{f}\ \left(\textrm{i.i.d.}\right).$ (21)
Note that the form of the distribution is not necessarily restricted to uniform one, where its variances scale as $O\left(d^{2}\right)$ (i.i.d.) is essential: This only affects the following $W_{2}$-KL equivalence up to the constant factor.
@@
$W_{2}^{2}\left(P:P^{\prime}\right)\sim\frac{2}{a^{2}}\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right].$ (25)
The proportionality constant explicitly reflects the scaling relation $h\left(d\right)=a\cdot d$, between the discretization length and the induced parameter uncertainty or its variation.
For Gaussian families, where the coordinates $\xi$ corresponds to the mean vector $\mu$, and $a=1$, this equivalence provides a direct physical intuition in terms of the Gaussian translations as discussed. In more general statistical submanifolds, the same structure should be interpreted as an information-geometric statement: the discretization-induced cost measured by 2-Wasserstein distance extracts the second-order geometric structure of the KL divergence in terms of Fisher metric at the leading order of $d\to 0$, which is independent of the detailed functional form of the distributions.

### v3

@@ II Concept and Derivation
Eq.(II) certainly exhibits that the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure. This result establishes a clear geometric interpretation of the UCN: The cost to project continuous Gaussian onto a discrete configurational support is given by the cumulative information-geometric distance for expected parallel translations within the Gaussian statistical manifold. These contributions are invisible to existing KL-based comparisons restricted to discrete statistical manifolds, but are naturally captured by the present transport-information-geometric framework.
In a realistic discrete system on a lattice, the following three features are unavoidable: (i) $d$ takes a nonzero positive value, (ii) the domain of the CDOS is bounded, and (iii) the domain is generally asymmetric. All of these features are dominated by the underlying lattice. Therefore, to address the effect of UCN in realistic systems, numerical approaches, such as systematic comparisons between Eq. (II) and Eq. (8) under various conditions, are essential for disentangling the individual contributions of these three lattice-induced effects.
@@
In the previous sections, we demonstrated that the discrepancy between continuous Gaussian distribution and its discretized counterpart, measured by the 2-Wasserstein distance with a quadratic cost aligned with Fisher metric, converges in the vanishing discretization limit $d\to 0$ to the expectation of the KL divergence between infinitesimally translated continuous Gaussian. Although these derivations was shown for Gaussian families, the underlying mathematical structure indicates its broader applicability. We here briefly address the conditions where such a correspondence can be generalized.
@@
where $f$ corresponds to the number of effective degrees of freedom relevant to the description of the system interested (and, subsequently, to the transport problem). Then we consider the problem to measure the difference between function $P\in S$ and its discretized counterpart $P^{\prime}$, based on $W_{2}$ and KL-divergence as discussed for Gaussian families. We assume that, in these coordinates, the Fisher metric (or its block-diagonal component associated with the transported degrees of freedom) is given by $f\times f$ positive-definite matrix of $\Omega$, which is taken as constant. More generally, our statement rely on the locally-constant character of $\Omega$ under infinitesimal variations in $\xi$. Then we define the quadratic cost function for the 2-Wasserstein distance $W_{2}\left(P:P^{\prime}\right)$ as
@@
$\rho\left(\delta\xi\right)=\textrm{Unif}\left[-\frac{h\left(d\right)}{2},\frac{h\left(d\right)}{2}\right]^{f}\ \left(\textrm{i.i.d.}\right),$ (21)
where the form of the distribution is not necessarily restricted to uniform one, where its variances scale as $O\left(d^{2}\right)$ (i.i.d.) is essential: This only affects the following $W_{2}$-KL equivalence up to the constant factor.
@@
$W_{2}^{2}\left(P:P^{\prime}\right)=\frac{2}{a^{2}}\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right].$ (25)
Remarks on the Generalization
The proportionality constant in Eq. (25) explicitly reflects the scaling relation $h\left(d\right)=a\cdot d$, between the discretization length and the induced parameter uncertainty or its variation. For Gaussian families, where the coordinates $\xi$ corresponds to the mean vector $\mu$ and $a=1$, this equivalence provides a direct physical intuition in terms of the Gaussian translations as discussed. In more general statistical submanifolds, the same structure should be interpreted as an information-geometric statement: the discretization-induced cost measured by 2-Wasserstein distance extracts the second-order geometric structure of the KL divergence in terms of Fisher metric at the leading order of $d\to 0$, which is independent of the detailed functional form of the distributions. We emphasize that the above leading-order correspondence relies on the assumption that the discretization-induced transport cost is dominated by local fluctuations: This condition can be naturally satisfied for distributions with sufficiently fast-decaying tails (including Gaussian families), where distant contributions to the quadratic transport cost are reasonably suppressed. In contrast, for distributions exhibiting heavy tails or slow decays, non-local contributions from the tail region may modify the scaling behavior of the $W_{2}$ distance in the limit $d\to 0$, potentially requiring corrections including the higher-order terms in $d$. The present analysis therefore characterizes a universal local correspondence, whose extension to heavy-tailed distributions is left for our future work.
Through the discussions about generalization, we now see that the extension of $\delta\xi$ to random vector is not merely a matter of average, but is structurally necessary: The trace structure appearing in the $W_{2}$ of Eq. (24) can only emerge in quadratic form of the KL divergence, if the outer product of $\delta\xi$ is effectively proportional to the identity matrix $I$, i.e., $\delta\xi\delta\xi^{\mathsf{T}}\propto d^{2}I_{f}$. Since the outer product of a single vector is necessarily rank one, while the required structure is full rank with $f$ degrees of freedom, an extension to a probabilistic ensemble of $\delta\xi$ is unavoidable for systems with $f\geq 2$.
Finally, we comment on the role of parametrization and invariance in the present Wasserstein-KL correspondence. While the KL divergence itself is invariant under reparametrizations on the statistical manifold, Eqs. (24)-(25) suggest that its expectation under $\rho$ depends on the trace of Fisher metric $\Omega$. This does not represent a contradiction, but can be reasonably understood from the following two aspects. (i) Taking the expectation $\mathbb{E}_{\rho}$ in Eq. (25) is an extrinsic operation living outside the statistical manifold, where the infinitesimal change in the selected parameter, $\delta\xi$, is linked to the discretization scale $d$ and subsequently extended to a random vector (the associated probability density $\rho$ typically reflects a modeling choice related to discretization or coarse-graining, depending on the problem under consideration). (ii) In general, when higher-order terms are present, the KL divergence expanded up to second order in $\delta\xi$, yielding a quadratic form, is explicitly governed by the Fisher metric $\Omega$ associated with the chosen parameter $\xi$. In such cases, once combined with the discretization scale and the extrinsic expectation over parameter variations, the resulting numerical value depends on the chosen parameter and Fisher metric. From these viewpoints, the explicit expression $d^{2}\mathrm{Tr}(\Omega)/12$ generally depends on the chosen parameter, yet the correspondence between the 2-Wasserstein transport cost and the expected KL divergence should be understood as an invariant structural relation. In particular, for the case of UCN in this work, since the Fisher metric is naturally identified through the isomorphism between the parameter $\mu$ on the statistical manifold and the transport coordinate, and the parameter variations scale linearly with the discretization length, the proportionality constant can be fixed. As a result, the correspondence yields an equality, Eq. (II), rather than a mere proportionality. In more general settings, different parameters and modeling choices for parameter variations lead to different numerical prefactors. Nevertheless, the universal correspondence $W_{2}^{2}\propto 2/a^{2}\cdot\mathbb{E}_{\rho}[D]$ is always preserved, expressing the robust emergence of second-order information-geometric structure from discretization-induced transport costs.